# Generative Medical Event Models Improve with Scale

### v1

@@ 4.1 Epic Cosmos
As described above, Cosmos is a collaboration for integrating and linking electronic health record (EHR) data between 310 healthcare organizations [^2] https://cosmos.epic.com/community in the United States, Canada, Lebanon, and Saudi Arabia as of August 2025 [6]. The Cosmos population from healthcare organizations in the United States is largely representative of census data. While some variation from national statistics has been noted, Cosmos has been shown to accurately reflect information and trends in other national databases (e.g. transplant registry, comorbidity patterns) [45, 46]. At the time of writing, most of the 121 published studies using Cosmos focus on describing the epidemiology or outcomes of specific conditions or services [47, 48, 49].
The current work is distinct insofar as most of our objectives are predictive rather than descriptive or associative—we predict diagnoses, acute exacerbation of chronic disease, and utilization outcomes. With that in mind, some of the conditions and outcomes we explored in this evaluation have been studied at a smaller scale in previous research using Cosmos data. These studies highlight challenges and opportunities. Varghese et al. [50] used Cosmos to demonstrate that the type of subtle differences CoMET captures impact risk of acute events among patients with T2DM. Patel et al. [51] demonstrated that Cosmos rates of systemic lupus erythematosus (SLE) based on structured data in diagnosis codes were lower than expected, and paradoxical mortality findings in [52] highlight the challenges of working with EHR data. Relatedly, Cosmos has been used for developing and validating risk scores and strata in other conditions, including cancer-associated thrombosis [53], vision impairment in multiple sclerosis [54], and perinatal cardiovascular events among patients with rare congenital diseases [55]. Several Cosmos publications have studied healthcare utilization outcomes in distinct subpopulations  [56, 57, 58]. Section 2.6 demonstrates how forecasting these types of metrics can be scaled with CoMET. While the present evaluation did not include the detailed phenotype modeling and validation included in some of these studies, CoMET data can readily be analyzed and fine tuned with greater specificity and logical complexity employed in other Cosmos studies. In this way, CoMET may inform and be informed by more traditional Cosmos research.
@@ 4.2 Scaling laws for LLMs
Research into the scaling behavior of machine learning models has fundamentally shaped the trajectory of large language models. A pivotal study by Kaplan et al. [35] showed that as we increase a model’s size, the amount of training data, and the compute budget, the performance improves following a smooth power-law trend—a predictable pattern that governs how model performance grows with more resources. Henighan et al. [59] extended this analysis beyond text to other domains like images, video, and multi-modal tasks, finding that larger transformer models consistently yield better predictive power across modalities as compute and model size grow, again following power-law improvement curves. A further demonstration of scale came from Brown et al. [60], who presented 175B parameter GPT-3 and found that simply making the model extremely large unlocked emergent capabilities; notably the ability to perform new tasks in a few-shot setting without any task-specific training. Subsequently, Hoffmann et al. [34] observed that many earlier large models were trained with too little data. They trained a 70B Chinchilla with four times more data than GPT-3, using the same compute budget, and found it outperformed much larger models. The frontier of scale was later pushed further by Chowdhery et al. [61] and OpenAI [62, 63], where the performance continued to improve log-linearly with scale. Studies of LLM scaling laws provide a guiding principle for our work, which asks whether similar predictable gains occur when we scale up models on more patient data.
@@ 4.3 Medical foundation models
Inspired by the success of foundation models in general domains, researchers have started developing large-scale models tailored to electronic health records. BEHRT [64] introduced a BERT-like transformer model trained on the longitudinal medical histories of 1.6 million patients. This approach yielded substantial improvements for disease prediction compared to prior state-of-the-art models. Med-BERT [65] was developed from 28 million patients, which further validated the power of large-scale pretraining on structured EHR data. CLMBR [13] introduced an autoregressive next-code predictor trained on 3.4 million patient records; the representations it learned boosted AUCROC across multiple downstream tasks, especially in low-data settings. MOTOR [14] trained a transformer-based model for time-to-event (TTE) prediction using 55 million patient records, demonstrating the transfer learning ability for the TTE foundation model.
Moving beyond encoder-only models like BERT, recent research has explored generative and autoregressive transformers that can explicitly model the sequence of events in a patient’s timeline. CEHR-GPT [66] was one of the first attempts to train GPT models on structured EHR data. It showed the synthetic data generated by the model effectively captures the intricate patterns present in EHR data. Introduced by Renc et al. [16], ETHOS is trained on tokenized event streams of patient health timelines and tasked with predicting the next set of events in a patient’s record. Interestingly, it does so in a zero-shot fashion, effectively learning a simulator of possible futures for a patient. In a similar vein, Event Stream GPT [39] provides tooling to convert complex, irregular medical event sequences into a format that a transformer can ingest, and methods to handle the causally ordered generation of events. TransformEHR [67] is an encoder–decoder transformer pretrained with a “visit masking” strategy: it masks out all the medical codes in some future visit and trains the model to reconstruct them from the preceding history.
@@
While these EHR-focused foundation models have shown encouraging results, they also highlight challenges. A recent comprehensive review by Wornow et al. [37] examined over 80 such models and found that many were trained on relatively narrow datasets (e.g., a single hospital’s records) and evaluated on surrogate tasks that may not translate to real clinical impact. EHRSHOT [68] and FoMoH [69] introduced new benchmark suites designed around patient timelines that extended beyond intensive care settings. These benchmarks emphasize robust, fair, and clinically meaningful evaluations. Instead of zero-shot approaches, both studies focused on smaller-scale models ($\sim\!100$M parameters) trained using a pretrain-then-finetune paradigm. Recently, Zhang et al. [17] conducted the first study on scaling laws for EHR foundation models. They investigated how model performance scales with size and data volume on smaller scales using MIMIC-IV [70], identifying consistent patterns such as power-law relationships between compute resources, model parameters, and clinical utility.
@@ 5.1 Data Pre-Processing
This pipeline yields a cohort whose records are dense enough to train a sequence model while minimizing noise from sporadic documentation and non-clinical artifacts. We applied very minimal filtering of patient records or input data to reflect the diversity of patients and the realities of real world healthcare documentation. For a full breakdown of our dataset, see Table 4.
@@ 5.2 Tokenization Details
Our tokenization method adapts a few key techniques from Renc et al. [16]. In general, medical events are placed in chronological order according to the instant at which they were documented, with some type-specific nuances noted below. The token vocabulary is defined a priori based on the individual events that are possible rather than frequency-based methods like byte-pair encoding. When possible we tokenized using ontologies with codes that directly capture hierarchical and categorical information, such as ICD-10-CM and Anatomical Therapeutic Chemical (ATC). All CoMET models have a vocabulary size of 7,105, which is summarized in Table 2.
@@
Table 2: Medical events included in CoMET’s vocabulary. The tokenizer also includes separation, padding, and unknown tokens.
@@ 5.2.2 Encounters
Encounters are book-ended with start- and end-encounter tokens—each with a denoted encounter type (e.g., “Emergency_Start”, “Emergency_End”). Each encounter header contains the start token, a department specialty token, and possible chief complaint tokens. Chief complaints in Cosmos consist of a name and an optional body location, which we separated into two unique consecutive tokens per chief complaint. When multiple encounters overlap, their encounter header and end-encounter tokens each appear at the proper instant in the timeline. In general, we do not enforce rules for overlapping or nested encounter-start or -end tokens; their placement in the patient sequence corresponds only to the time at which the event happened, with the tokens for a given encounter header always appearing consecutively.
@@ 5.2.3 Diagnoses
Diagnosis events are represented by the associated ICD-10-CM code. ICD-10-CM codes were split up into three tokens by category and sub-category (first 3 characters), specific details (characters 4-5), and additional details and extensions (characters 6-7). Diagnoses are represented in CoMET patient trajectories by 1-3 tokens depending on the specificity of the documentation in the patient’s chart. Because diagnosis events in Cosmos only have documentation resolution at the day level, they are placed at the very beginning of the encounter after the encounter header or at midnight of their documented date, whichever comes later. Diagnoses with the same date are sequenced in random order.
@@ 5.2.4 Labs
Our dataset contains the 1,000 most common numeric lab tests from Cosmos (representing $99\%$ of all numeric lab test results). Each lab result is represented by two consecutive tokens which identify the test performed and the quantile of the numeric result, following ETHOS [16]. Lab components are represented by Logical Observation Identifiers Names and Codes (LOINC®) [^3] LOINC® is a registered trademark of Regenstrief Institute, Inc. codes, and numeric results are stratified by LOINC and measurement unit. For each LOINC-unit pair we partitioned values into 10 equal-frequency bins and mapped them to generic tokens. Compared to uniform-width binning, quantiles better accommodate non-linear mappings between heterogeneous units and better balance token frequencies, though they can compress rare but clinically meaningful extremes. The lab-result token pairs are placed in the sequence at the instance of collection. This ensures the sequence mirrors the patient’s evolving clinical state but calls for careful consideration to ensure that only information available at the evaluation time is used.
@@ 5.2.5 Medications
Medication orders are represented by an ATC code without other data. ATC codes were split up into three sets of tokens by anatomical group and therapeutic subgroup (characters 1-3), pharmacological and chemical subgroup (characters 4-5), and chemical substance (characters 6-7). Medication orders are represented in CoMET patient trajectories using three consecutive tokens to represent the full ATC code. They are placed in the sequence at the instant of the order.
@@ 5.2.8 Excluded Data
CoMET is limited to structured data and does not include data such as clinical notes, images, or free-text results from diagnostic procedures. Additional structured data, such as vitals, allergies, and medication administrations, were not included at this time.
@@ 5.3 Model Training Details
All CoMET models are built on the Qwen2 architecture. An overview of the architectural hyperparameters is shown in Table 3. All variants are trained with a context window of 8,192 tokens.
@@
CoMET‑S 62M 6 768 12 3072
CoMET‑M 119M 12 768 12 3072
CoMET‑L 1B 16 2048 32 8192
Table 3: Summary of CoMET model configurations, including Small (S), Medium (M), and Large (L) model sizes. Abbreviation: MLP = multi-layer perceptron
5.3.1 CoMET training and scaling laws
CoMET models are built on the Qwen2 architecture [18]. Qwen2 incorporates pre-layer normalization, SwiGLU activations, rotary positional embeddings, and grouped-query attention, all of which improve training time and stability. CoMET models were trained with cross-entropy loss and standard hyperparameters, slightly adjusting them as determined by experimentation to account for possible differences in training transformers on medical events rather than natural language. Batch size was fixed at 512 sequences, and input sequences were densely packed in order to fully use the context window during training, with only a separation token marking different patients since our experiments demonstrated this was sufficient for training. Following Hoffmann et al. [34], we employed a 10x learning rate decay with cosine schedule and used AdamW as our optimizer.
@@ 5.4 Benchmarking with Task-Specific Supervised Models
To contextualize the performance of CoMET, we implemented three baseline approaches representing distinct modeling paradigms. These included (1) linear and logistic regression models, representing classical linear approaches; (2) gradient-boosted decision trees, a strong non-linear method widely used for structured data; and (3) supervised transformers trained from a random initialization, a flexible deep learning architecture that retains the temporal information of each patient record. Each baseline model was trained independently for its corresponding downstream task and evaluated using the same datasets and procedures as those applied to the CoMET foundational models.
@@
For each downstream task, we applied the evaluation’s labeling logic to the CoMET training set to construct supervised examples. We used all supervised examples available in the CoMET training dataset up to a maximum of 5 million patient histories per task to fit the baselines. These samples were grouped by patient id and stratified by the task label where applicable. We reserved an additional 625,000 stratified rows when available or $10\%$ of the available training rows as a validation set for early stopping and hyper-parameter selection, leaving the designated CoMET development and test splits untouched for final evaluation.
@@
With the exception of the supervised transformer baselines, input prompts were converted into a bag-of-words (BoW) count vector over the full CoMET vocabulary (Table 2). Columns were normalized such that the maximum absolute value of each feature in the training set was 1.0 before fitting the linear and logistic models; XGBoost consumed raw counts. Preliminary experiments with inverse-frequency class-weighting showed no material AUCROC gain on these large samples, so we report unweighted results for simplicity.
@@
We used a gradient-boosted decision trees classifier or regressor with n_estimators=10,000=10{,}000, max_depth=6=6, learning_rate=0.1=0.1, and subsample=0.8=0.8. Training employed early stopping with a patience of $100$ boosting rounds. Other hyper-parameters followed XGBoost 1.5.2 library defaults unless specified.
@@
We trained the closest task-specific supervised counterpart to CoMET-M by using the same architecture as CoMET-M (Table 3) and the same tokenized input prompts. Unlike the bag-of-words inputs, this baseline preserves model architecture and the temporal sequence information in the original prompts. We attached a task-specific classification or regression head to the model and optimized with AdamW (weight decay 0.01), using a linear-decay schedule with warmup (500 warmup steps) and a peak learning rate of $1\times 10^{-4}$. Training ran for 10,00010{,}000 update steps (global batch size $128$; 1.281.28M samples). We applied gradient-norm clipping at $1.0$ and dropout 0.0.

### v3

@@ 4.1 Epic Cosmos
As described above, Cosmos is a collaboration for integrating and linking electronic health record (EHR) data among 310 healthcare organizations [^2] https://cosmos.epic.com/community in the United States, Canada, Lebanon, and Saudi Arabia as of August 2025 [6]. The Cosmos population from healthcare organizations in the United States is largely representative of U.S. census data. While some variation from national statistics has been noted, Cosmos has been shown to accurately reflect information and trends in other national databases (e.g. transplant registry, comorbidity patterns) [45, 46]. At the time of writing, most of the 121 published studies using Cosmos focus on describing the epidemiology or outcomes of specific conditions or services [47, 48, 49].
The current work is distinct insofar as most of our objectives are predictive rather than descriptive or associative—we predict diagnoses, acute exacerbations of chronic disease, and utilization outcomes. With that in mind, some of the conditions and outcomes we explored in this evaluation have been studied at a smaller scale in previous research using Cosmos data. These studies highlight challenges and opportunities. Varghese et al. [50] used Cosmos to demonstrate that the types of subtle differences Curiosity captures impact risk of acute events among patients with T2DM. Patel et al. [51] demonstrated that Cosmos rates of systemic lupus erythematosus (SLE) based on structured data in diagnosis codes were lower than expected, and paradoxical mortality findings in [52] highlight the challenges of working with EHR data. Relatedly, Cosmos has been used for developing and validating risk scores and strata in other conditions, including cancer-associated thrombosis [53], vision impairment in multiple sclerosis [54], and perinatal cardiovascular events among patients with rare congenital diseases [55]. Several Cosmos publications have studied healthcare utilization outcomes in distinct subpopulations  [56, 57, 58]. Section 2.6 demonstrates how forecasting these types of metrics can be scaled with Curiosity. While the present evaluation did not include the detailed phenotype modeling and validation present in some of these studies, Curiosity data can readily be analyzed and fine-tuned with greater specificity and logical complexity employed in other Cosmos studies. In this way, Curiosity may inform and be informed by more traditional Cosmos research.
@@ 4.2 Scaling laws for LLMs
Research into the scaling behavior of machine learning models has fundamentally shaped the trajectory of large language models. A pivotal study by Kaplan et al. [35] showed that as we increase a model’s size, the amount of training data, and the compute budget, the performance improves following a smooth power-law trend—a predictable pattern that governs how model performance grows with more resources. Henighan et al. [59] extended this analysis beyond text to other domains like images, video, and multimodal tasks, finding that larger transformer models consistently yield better predictive power across modalities as compute and model size grow, again following power-law improvement curves. A further demonstration of scale came from Brown et al. [60], who presented 175B-parameter GPT-3 and found that simply making the model extremely large unlocked emergent capabilities; notably the ability to perform new tasks in a few-shot setting without any task-specific training. Subsequently, Hoffmann et al. [34] observed that many earlier large models were trained with too little data. They trained a 70B-parameter Chinchilla with four times more data than GPT-3, using the same compute budget, and found it outperformed much larger models. The frontier of scale was later pushed further by Chowdhery et al. [61] and OpenAI [62, 63], where performance continued to improve log-linearly with scale. Studies of LLM scaling laws provide a guiding principle for our work, which asks whether similar predictable gains occur when we scale up models on more patient data.
@@ 4.3 Medical foundation models
Inspired by the success of foundation models in general domains, researchers have started developing large-scale models tailored to electronic health records. BEHRT [64] introduced a BERT-like transformer model trained on the longitudinal medical histories of 1.6 million patients. This approach yielded substantial improvements for disease prediction compared to prior state-of-the-art models. Med-BERT [65] was developed from 28 million patients, which further validated the power of large-scale pretraining on structured EHR data. CLMBR [13] introduced an autoregressive next-code predictor trained on 3.4 million patient records; the learned representations improved AUCROC across multiple downstream tasks, especially in low-data settings. MOTOR [14] trained a transformer-based model for time-to-event (TTE) prediction using 55 million patient records, demonstrating the transfer learning ability for the TTE foundation model.
Moving beyond encoder-only models like BERT, recent research has explored generative and autoregressive transformers that can explicitly model the sequence of events in a patient’s timeline. CEHR-GPT [66] was one of the first attempts to train GPT models on structured EHR data. It showed that the synthetic data generated by the model effectively captures the intricate patterns present in EHR data. Introduced by Renc et al. [16], ETHOS is trained on tokenized event streams of patient health timelines and tasked with predicting the next set of events in a patient’s record. Interestingly, it does so in a zero-shot fashion, effectively learning a simulator of possible futures for a patient. In a similar vein, Event Stream GPT [39] provides tooling to convert complex, irregular medical event sequences into a format that a transformer can ingest, and methods to handle the causally ordered generation of events. TransformEHR [67] is an encoder–decoder transformer pretrained with a “visit masking” strategy: it masks out all the medical codes in some future visit and trains the model to reconstruct them from the preceding history.
@@
While these EHR-focused foundation models have shown encouraging results, they also highlight challenges. A recent comprehensive review by Wornow et al. [37] examined over 80 such models and found that many were trained on relatively narrow datasets (e.g., a single hospital’s records) and evaluated on surrogate tasks that may not translate to real clinical impact. EHRSHOT [68] and FoMoH [69] introduced new benchmark suites designed around patient timelines that extend beyond intensive care settings. These benchmarks emphasize robust, fair, and clinically meaningful evaluations. Instead of zero-shot approaches, both studies focused on smaller-scale models ($\sim\!100$M parameters) trained using a pretrain-then-finetune paradigm. Recently, Zhang et al. [17] conducted the first study on scaling laws for EHR foundation models. They investigated how model performance scales with size and data volume at smaller scales using MIMIC-IV [70], identifying consistent patterns such as power-law relationships between compute resources, model parameters, and clinical utility.
@@ 5.1 Data Pre-Processing
This pipeline yields a cohort whose records are dense enough to train a sequence model while minimizing noise from sporadic documentation and non-clinical artifacts. We applied minimal filtering of patient records or input data to reflect the diversity of patients and the realities of real-world healthcare documentation. For a full breakdown of our dataset, see Table 4.
@@ 5.2 Tokenization Details
Our tokenization method adapts a few key techniques from Renc et al. [16]. In general, medical events are placed in chronological order according to the instant at which they were documented, with some type-specific nuances noted below. The token vocabulary is defined a priori based on the individual events that are possible rather than frequency-based methods like byte-pair encoding. When possible we tokenized using ontologies with codes that directly capture hierarchical and categorical information, such as ICD-10-CM and Anatomical Therapeutic Chemical (ATC). All Curiosity models have a vocabulary size of 7,105, which is summarized in Table 2.
@@
Table 2: Medical events included in Curiosity’s vocabulary. The tokenizer also includes separation, padding, and unknown tokens.
@@ 5.2.2 Encounters
Encounters are bookended with start- and end-encounter tokens—each with a denoted encounter type (e.g., “Emergency_Start”, “Emergency_End”). Each encounter header contains the start token, a department specialty token, and possible chief complaint tokens. Chief complaints in Cosmos consist of a name and an optional body location, which we separated into two unique consecutive tokens per chief complaint. When multiple encounters overlap, their encounter header and end-encounter tokens each appear at the proper instant in the timeline. In general, we do not enforce rules for overlapping or nested encounter-start or -end tokens; their placement in the patient sequence corresponds only to the time at which the event happened, with the tokens for a given encounter header always appearing consecutively.
@@ 5.2.3 Diagnoses
Diagnosis events are represented by the associated ICD-10-CM code. ICD-10-CM codes were split up into three tokens by category and sub-category (first 3 characters), specific details (characters 4-5), and additional details and extensions (characters 6-7). Diagnoses are represented in Curiosity patient trajectories by 1-3 tokens depending on the specificity of the documentation in the patient’s chart. Because diagnosis events in Cosmos only have documentation resolution at the day level, they are placed at the very beginning of the encounter after the encounter header or at midnight of their documented date, whichever comes later. Diagnoses with the same date are sequenced in random order.
@@ 5.2.4 Labs
Our dataset contains the 1,000 most common numeric lab tests from Cosmos (representing $99\%$ of all numeric lab test results). Each lab result is represented by two consecutive tokens which identify the test performed and the quantile of the numeric result, following ETHOS [16]. Lab components are represented by Logical Observation Identifiers Names and Codes (LOINC®) [^3] LOINC® is a registered trademark of Regenstrief Institute, Inc. codes, and numeric results are stratified by LOINC and measurement unit. For each LOINC-unit pair we partitioned values into 10 equal-frequency bins and mapped them to generic tokens. Compared to uniform-width binning, quantiles better accommodate non-linear mappings between heterogeneous units and better balance token frequencies, though they can compress rare but clinically meaningful extremes. The lab-result token pairs are placed in the sequence at the instant of collection. This ensures the sequence mirrors the patient’s evolving clinical state but requires careful consideration to ensure that only information available at evaluation time is used.
@@ 5.2.5 Medications
Medication orders are represented by an ATC code without other data. ATC codes were split up into three sets of tokens by anatomical group and therapeutic subgroup (characters 1-3), pharmacological and chemical subgroup (characters 4-5), and chemical substance (characters 6-7). Medication orders are represented in Curiosity patient trajectories using three consecutive tokens to represent the full ATC code. They are placed in the sequence at the instant of the order.
@@ 5.2.8 Excluded Data
Curiosity is limited to structured data and does not include data such as clinical notes, images, or free-text results from diagnostic procedures. Additional structured data, such as vitals, allergies, and medication administrations, were not included at this time.
@@ 5.3 Model Training Details
All Curiosity models are built on the Qwen2 architecture. An overview of the architectural hyperparameters is shown in Table 3. All variants are trained with a context window of 8,192 tokens.
@@
Curiosity-S 62M 6 768 12 3072
Curiosity-M 119M 12 768 12 3072
Curiosity-L 1B 16 2048 32 8192
Table 3: Summary of Curiosity model configurations, including Small (S), Medium (M), and Large (L) model sizes. Abbreviation: MLP = multi-layer perceptron
5.3.1 Curiosity training and scaling laws
Curiosity models are built on the Qwen2 architecture [18]. Qwen2 incorporates pre-layer normalization, SwiGLU activations, rotary positional embeddings, and grouped-query attention, all of which improve training time and stability. Curiosity models were trained with cross-entropy loss and standard hyperparameters, slightly adjusting them as determined by experimentation to account for possible differences in training transformers on medical events rather than natural language. Batch size was fixed at 512 sequences, and input sequences were densely packed in order to fully use the context window during training, with only a separation token marking different patients since our experiments demonstrated this was sufficient for training. Following Hoffmann et al. [34], we employed a 10x learning rate decay with cosine schedule and used AdamW as our optimizer.
@@ 5.4 Benchmarking with Task-Specific Supervised Models
To contextualize the performance of Curiosity, we implemented three baseline approaches representing distinct modeling paradigms. These included (1) linear and logistic regression models, representing classical linear approaches; (2) gradient-boosted decision trees, a strong non-linear method widely used for structured data; and (3) supervised transformers trained from a random initialization, a flexible deep learning architecture that retains the temporal information of each patient record. Each baseline model was trained independently for its corresponding downstream task and evaluated using the same datasets and procedures as those applied to the Curiosity foundational models.
@@
For each downstream task, we applied the evaluation’s labeling logic to the Curiosity training set to construct supervised examples. We used all supervised examples available in the Curiosity training dataset up to a maximum of 5 million patient histories per task to fit the baselines. These samples were grouped by patient id and stratified by the task label where applicable. We reserved an additional 625,000 stratified rows when available or $10\%$ of the available training rows as a validation set for early stopping and hyper-parameter selection, leaving the designated Curiosity development and test splits untouched for final evaluation.
@@
With the exception of the supervised transformer baselines, input prompts were converted into a bag-of-words (BoW) count vector over the full Curiosity vocabulary (Table 2). Columns were normalized such that the maximum absolute value of each feature in the training set was 1.0 before fitting the linear and logistic models; XGBoost consumed raw counts. Preliminary experiments with inverse-frequency class-weighting showed no material AUCROC gain on these large samples, so we report unweighted results for simplicity.
@@
We used a gradient-boosted decision tree classifier or regressor with n_estimators=10,000=10{,}000, max_depth=6=6, learning_rate=0.1=0.1, and subsample=0.8=0.8. Training employed early stopping with a patience of $100$ boosting rounds. Other hyper-parameters followed XGBoost 1.5.2 library defaults unless specified.
@@
We trained the closest task-specific supervised counterpart to Curiosity-M by using the same architecture as Curiosity-M (Table 3) and the same tokenized input prompts. Unlike the bag-of-words inputs, this baseline preserves model architecture and the temporal sequence information in the original prompts. We attached a task-specific classification or regression head to the model and optimized with AdamW (weight decay 0.01), using a linear-decay schedule with warmup (500 warmup steps) and a peak learning rate of $1\times 10^{-4}$. Training ran for 10,00010{,}000 update steps (global batch size $128$; 1.281.28M samples). We applied gradient-norm clipping at $1.0$ and dropout 0.0.